OpenStudy (anonymous):

f(x)=x^2+6x+25 find the zeroes

OpenStudy (gorv):

we need to find factor of it

OpenStudy (gorv):

break 6 in 2 parts in such a way that sum of two broken part= 6 and product of those 2 part= 25

OpenStudy (mathstudent55):

Set the polynomial equal to zero. Then try factoring the polynomial. If that doesn't work, use the quadratic formula.

OpenStudy (anonymous):

how would i use he quadratic formula ?

OpenStudy (anonymous):

i figured it out, thanks.

OpenStudy (gorv):

u know what is quadratic formula??

OpenStudy (mathstudent55):

\(x = \dfrac{-b \pm \sqrt{b^2 - 4ac}}{2a} \) where a = 1; b = 6; c = 25